10g of the gas as 17oc are compressed adiabatically and reversibly from 8 to 5 liters.

i. Calculate the final temperature

ii. Calculate the work done on the gas.

iii. What are U and H for the process (Cp = 3/2R and R = 8.314).


1
Expert's answer
2021-06-28T08:10:38-0400

"\frac{V_1}{T_1}=\frac{V_2}{T_2}"


"\frac{8}{290}=\frac{5}{T_2}"


"T_2=181.25k"


Work done = 290 - 181.25=108.75